OpenStudy (anonymous):

Help with chemistry lab please asap? I need to determine the formula of a sample of copper sulfate hydrate.? Crucible Crucible & hydrate Dehydrate 20.500 g 21.255 g 2 0.983 g This is what I did so far: An empty crucible has a mass of 20.500g. A sample of copper sulfate hydrate is added to the crucible, making the total mass of the crucible and hydrate 21.255 g. Subtraction of the mass of the empty crucible from the mass of the hydrate and crucible gives the mass of the hydrate. (21.255 g- 20.500 g = 0.755 g hydrate).

OpenStudy (anonymous):

continued: Step 2: Subtract the mass of the dehydrated compound from the mass of the hydrate to determine the mass of water that was present in the hydrate. 20.983 g - 0.755 g = 20.228 g H2O present in the sample of hydrate

OpenStudy (anonymous):

please show work. thanks

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

@chmvijay

OpenStudy (anonymous):

I only have the text version of the lab, which is this:

OpenStudy (anonymous):

Molar Mass of Compounds: Determining the Formula of a Hydrate—Text Version When hydrates are heated, the water of hydration evaporates and is released as vapor. The remaining solid is known as the anhydrous salt, which is the compound same compound without the water molecules trapped inside. Purpose: To determine the formula of a hydrate. Materials: • crucible •Bunsen burner •balance •CuSO4 hydrate Procedure: 1.Measure the mass of the clean, empty crucible, record the mass. 2.Add one or two scoops of the hydrate to the crucible, record the mass. 3.Heat the crucible and hydrate above a Bunsen burner for at least ten minutes to make sure that all of the water evaporates. 4.Let the crucible cool and then mass the crucible with dehydrated solid inside, record the mass. 5.Calculate the formula of the hydrate.

OpenStudy (anonymous):

I basically need to know how to find the empiricial formula

OpenStudy (aaronq):

convert the masses (of the dehydrated compound and the water) you found to moles and compare them.

OpenStudy (anonymous):

I know that, but idk how

OpenStudy (aaronq):

have you found the moles?

OpenStudy (anonymous):

I'm not good with moles convertion

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I need someone to explain it to me

OpenStudy (aaronq):

use this formula: \(n=\dfrac{m}{M}\rightarrow n_{CuSO_4}=\dfrac{0.983g}{159.62g/mol} \) do the same for the water

OpenStudy (aaronq):

n=moles, m=mass, M=molar mass

OpenStudy (anonymous):

idk what the molas mass is tho

OpenStudy (aaronq):

you can find it on a periodic table, just add up the molar masses of each individual atom in the compound

OpenStudy (anonymous):

I don't really know how to do that

OpenStudy (aaronq):

heres a periodic table http://ptable.com the molar masses are given at the bottom of each square for that element.

OpenStudy (anonymous):

so I add up the one from C, S, and O and it will be one number?

OpenStudy (anonymous):

ones*

OpenStudy (aaronq):

i already did the CuSO4 one. (i wrote it above), you can do the one for water

OpenStudy (anonymous):

I mean Cu

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

but how did u get 2 numbers

OpenStudy (anonymous):

wait, so for H its 1.008 and for O its 15.999

OpenStudy (aaronq):

the number at the top was the mass of the dehydrated substance, which is \(CuSO_4\). The hydrate is \(CuSO_4xH_2O\) so now you're finding how many moles of water there are

OpenStudy (aaronq):

okay, but there are 2 hydrogens, so you need to add that 2 times

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

1.008 * 2= 2.016

OpenStudy (aaronq):

+ the oxygen

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

18.015

OpenStudy (aaronq):

good stuff. now, can you find the mass of the water?

OpenStudy (anonymous):

does 18.015 go on the top of the fraction

OpenStudy (aaronq):

it goes where the uppercase M is, so at the bottom.

OpenStudy (anonymous):

ok how do i find the mass of H2o

OpenStudy (aaronq):

you can find it with the data given in the question Crucible Crucible & hydrate Dehydrate 20.500 g 21.255 g 2 0.983 g mass of hydrate - mass of dehydrate = mass of water

OpenStudy (anonymous):

wait hello we arent done??

OpenStudy (anonymous):

@aaronq now what.. i need the formula

OpenStudy (aaronq):

have you found the mass of the water?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

0.272 g

OpenStudy (aaronq):

now find the moles by diving by the molar mass

OpenStudy (anonymous):

so its n H2O= 0.272 g/18.015 g/mol

OpenStudy (aaronq):

okay good. so what are the moles of each

OpenStudy (anonymous):

So I divide? do i also do it to CuSO4?

OpenStudy (aaronq):

yes

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

for water, its a long decimal

OpenStudy (anonymous):

0.0150985...

OpenStudy (aaronq):

0.015 is fine

OpenStudy (anonymous):

the next one is 0.0061..

OpenStudy (anonymous):

0.006158...

OpenStudy (aaronq):

okay. so now that we know the amount of moles (which correlates to the number of atoms) we can compare the 2 to see in what ratio the hydrate is made of. so divide the moles of water by the moles of the CuSO4: \(\dfrac{n_{H_2O}}{n_{CUSO_4}}\)

OpenStudy (anonymous):

wait hold on, i forgot to say that in the data table we only have the mass of the Crucible and hydrate together, but i also calulated the mass of ONLY the hydrate, which is 0.755 g. do we need to use that?

OpenStudy (aaronq):

nope. we only need the mass of the dehydrate and the water

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

its 0.05514...

OpenStudy (aaronq):

you divided incorrectly: 0.0150985/0.0061= 2.47

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

because i didnt do the whole decimal

OpenStudy (anonymous):

wait but the full decimal for CUSO4 isn't 0.0061

OpenStudy (anonymous):

its 0.006158 and it continues.. i dnt remember the rest

OpenStudy (anonymous):

i can divide again

OpenStudy (anonymous):

do we need to use the wholeee decimal to divide?

OpenStudy (aaronq):

i'm checking your calculations and the mass of the water you found is incorrect. 0.983 -( 21.255-20.500) = 0.228 g the moles of water are 0.0126666 the moles of the dehydrate are 0.0061583 so the ratio is: 0.0126666/0.0061583=2.05683386

OpenStudy (aaronq):

no you don't, but you placed the numbers in the wrong places.

OpenStudy (anonymous):

oh

OpenStudy (aaronq):

so, you have 2 water molecules for every molecules of copper sulfate, so the formula is \(CuSO_4*2H_2O\)

OpenStudy (anonymous):

so for moles of water is this right? N= m/M  n H2O= 0.272 g/18.015 g/mol= 0.0126666

OpenStudy (aaronq):

yes that's right

OpenStudy (anonymous):

ok

OpenStudy (aaronq):

do you understand that last part?

OpenStudy (anonymous):

yes

OpenStudy (aaronq):

okay cool.

OpenStudy (anonymous):

oh my god u are a lifesaver . thanks

OpenStudy (aaronq):

no problem !